Project

General

Profile

2017-12-05-wedge

Present:
  • Pavel
  • Dave Neuffer
  • Tanaz
  • Chris Rogers
Chris outlined the history as-of 2 pm today
  • discussion of attempt to bring back SSD
    Tanaz said
  • M2D ramped to 50 A
  • ECE ramping to 50 A
    By the end of the meeting ECE was held at 50 A.
Empty LiH Settings taken to date (all solenoid mode settings):
  • 2017-02-6 (140 MeV/c M2D ON)
  • 2016-04 1.2 (140 MeV/c M2D OFF)
  • 2016-04 2.3 (200 MeV/c M2D OFF)
  • 2016-04 2.4 (240 MeV/c M2D OFF)

We believe plan is to attempt to ramp to a M2D off setting tonight; and ramp to a M2D on setting tomorrow

Tanaz showed slides. Indicate that there is no longitudinal density increase in 2016/05 1 setting (140 MeV/c, flip mode, M2D OFF). Indication is that we would like M2D ON for the wedge programme.

Slide.pdf - Tanaz slides showing wedge using 2016.05 1 (2.01 MB) Rogers, Chris, 05 December 2017 16:47